5. Suppose that a random sample of 25 mining caps was tested from a population of 300 mining caps and 20 exploded properly. Construct a 95% confidence interval for the proportion of mining caps that will explode properly.
1
Expert's answer
2020-10-06T18:05:57-0400

"\text{The confidence interval for the }\\
\text{proportion can be calculated as follows:}\\
 \hat p ±Z_{\frac{α}{2}}(\sqrt {\frac{\hat p (1-\hat p )}{n}}),\\
 \hat p =\frac{20}{25}=0.8,\\
 Z_{\frac{α}{2}}=Z_{\frac{0.05}{2}}=Z_{0.025}=1.96,\\
 \text{the lower limit }= 0.8-1.96(\sqrt {\frac{0.8 (1-0.8 )}{25}})\\
≈0.6432,\\
 \text{the upper limit }= 0.8+1.96(\sqrt {\frac{0.8 (1-0.8 )}{25}})\\
≈0.9568,\\
 \text{so the confidence interval is}\\
 0.6432\leq \ p \leq 0.9568"